# On a class of nonlinear BGK-type kinetic equations with density dependent collision rates

### v2

@@ 4.1 $L^{2}$-hypocoercivity à la Dolbeault-Mouhot-Schmeiser
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\int_{\mathbb{T}^{d}}\int_{\mathbb{R}^{d}}\frac{(f_{\varepsilon}-\mathcal{M})^{2}}{\mathcal{M}}\,\mathrm{d}v\,\mathrm{d}x$ $=-\frac{1}{\varepsilon^{2}}\int_{\mathbb{T}^{d}}\rho_{\varepsilon}^{\alpha}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left(g_{\varepsilon}^{2}(u)\frac{\mathcal{M}(v)}{\mathcal{M}(u)}+g_{\varepsilon}^{2}(v)\frac{\mathcal{M}(v)}{\mathcal{M}(u)}-g_{\varepsilon}(v)g_{\varepsilon}(u)\right)\,\mathrm{d}u\,\mathrm{d}v\,\mathrm{d}x$
$=-\frac{2}{\varepsilon^{2}}\int_{\mathbb{T}^{d}}\rho_{\varepsilon}^{\alpha}\int_{\mathbb{R}^{d}}\frac{g_{\varepsilon}^{2}}{\mathcal{M}}\,\mathrm{d}v$
@@
Now using Lemma 4.1 we estimate the above quantity further.
@@
$\qquad+\frac{1}{\varepsilon^{2}}\|\rho_{\varepsilon}^{\alpha}\|_{L^{\infty}}\left(\frac{\|\rho_{\varepsilon}^{\alpha}\|_{L^{\infty}}}{\varepsilon}\|j_{\varepsilon}\|_{L^{2}_{x}}^{2}+\frac{\varepsilon}{4\|\rho_{\varepsilon}^{\alpha}\|_{L^{\infty}}}\|\rho_{\varepsilon}-\bar{\rho_{\varepsilon}}\|_{L^{2}_{x}}^{2}\right)-\frac{1}{\varepsilon}\|\rho_{\varepsilon}-\bar{\rho_{\varepsilon}}\|_{L^{2}_{x}}^{2}$
$\leq-\frac{1}{2\varepsilon}\|\rho_{\varepsilon}-\bar{\rho_{\varepsilon}}\|_{L^{2}_{x}}^{2}+\frac{1}{\varepsilon}\|E_{\varepsilon}\|_{L^{2}_{x}}^{2}+C\|\rho_{\varepsilon}^{\alpha}\|_{L^{\infty}}^{2}\frac{1}{\varepsilon^{3}}\|j_{\varepsilon}\|_{L^{2}_{x}}^{2}$
$\leq-\frac{1}{2\varepsilon}\|\rho_{\varepsilon}-\bar{\rho_{\varepsilon}}\|_{L^{2}_{x}}^{2}+\frac{C}{\varepsilon^{3}}(1+\|\rho_{\varepsilon}^{\alpha}\|_{L^{\infty}}^{2})\|(I-\Pi)f_{\varepsilon}\|^{2}_{L^{2}_{x,v}(\mathcal{M}^{-1})},$ (4.5)
for some constant $C$ not depending on $\|\rho_{\varepsilon}\|_{L^{\infty}}$ or $\varepsilon$. Above, to get to the last line we used Lemma 2.4 with $\beta=\rho_{\varepsilon}$.
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\left(\|f_{\varepsilon}-\mathcal{M}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}+\frac{\varepsilon}{CA^{\alpha}(1+A^{2\alpha})}\int_{\mathbb{T}^{d}}j_{\varepsilon}\cdot(I-\Delta_{x})^{-1}\nabla_{x}\rho_{\varepsilon}\,\mathrm{d}x\right)\\ \leq-\frac{1}{\varepsilon^{2}}\frac{1}{A^{\alpha}}\|(I-\Pi)f_{\varepsilon}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}-\frac{1}{4CA^{\alpha}(1+A^{2\alpha})}\|\rho_{\varepsilon}-\bar{\rho_{\varepsilon}}\|_{L^{2}_{x}}^{2}.$ (4.6)
@@
$\frac{\varepsilon}{CA^{\alpha}(1+A^{2\alpha})}\int_{\mathbb{T}^{d}}j_{\varepsilon}\cdot(I-\Delta_{x})^{-1}\nabla_{x}\rho_{\varepsilon}\,\mathrm{d}x$ $\leq\frac{\varepsilon}{CA^{\alpha}(1+A^{2\alpha})}\|j_{\varepsilon}\|_{L^{2}_{x}}\|(I-\Delta_{x})^{-1}\nabla_{x}\rho_{\varepsilon}\|_{L^{2}_{x}}$
$\leq\frac{\varepsilon}{CA^{\alpha}(1+A^{2\alpha})}\|f_{\varepsilon}-\mathcal{M}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}\leq\frac{1}{2}\|f_{\varepsilon}-\mathcal{M}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})},$
with the second inequality following from Lemma 2.4 with $\beta=1$. Consequently for $\varepsilon$ sufficiently small, we obtain
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\Big{(}$ $\|f_{\varepsilon}-\mathcal{M}\|^{2}_{L^{2}(\mathcal{M}^{-1})}+\frac{\varepsilon}{CA^{\alpha}(1+A^{2\alpha})}\int_{\mathbb{T}^{d}}j_{\varepsilon}\cdot(I-\Delta_{x})^{-1}\nabla_{x}\rho_{\varepsilon}\,\mathrm{d}x\Big{)}$
$\leq-\frac{1}{4CA^{\alpha}(1+A^{2\alpha})}\bigl{(}\|(I-\Pi)f_{\varepsilon}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}-\|\rho_{\varepsilon}-\bar{\rho_{\varepsilon}}\|_{L^{2}_{x}}^{2}\bigr{)}$
$=-\frac{1}{4CA^{\alpha}(1+A^{2\alpha})}\|f_{\varepsilon}-\mathcal{M}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}$
$\leq-\frac{1}{8CA^{\alpha}(1+A^{2\alpha})}\left(\|f_{\varepsilon}-\mathcal{M}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}+\frac{\varepsilon}{CA^{\alpha}(1+A^{2\alpha})}\int_{\mathbb{T}^{d}}j_{\varepsilon}\cdot(I-\Delta_{x})^{-1}\nabla_{x}\rho_{\varepsilon}\,\mathrm{d}x\right).$
@@
$\|f_{\varepsilon}(t)-\mathcal{M}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}+\frac{\varepsilon}{CA^{\alpha}(1+A^{2\alpha})}\int_{\mathbb{T}^{d}}j_{\varepsilon}(t)\cdot(I-\Delta_{x})^{-1}\nabla_{x}\rho_{\varepsilon}(t)\,\mathrm{d}x\\ \leq e^{-\gamma t}\left(\|f_{\varepsilon}(0)-\mathcal{M}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}+\frac{\varepsilon}{CA^{\alpha}(1+A^{2\alpha})}\int_{\mathbb{T}^{d}}j_{\varepsilon}(0)\cdot(I-\Delta_{x})^{-1}\nabla_{x}\rho_{\varepsilon}(0)\,\mathrm{d}x\right),$
@@ 5.1 Long-time asymptotics
Notice that Equation (1.3) is uniformly parabolic away from zero and +$\infty$+ ∞, so the classical theory applies for initial data uniformly bounded from above and below. We then have well-posedness and a strong maximum principle and Lemma 5.1 follows. We refer the reader to [36] for a review of the theory of the nonlinear diffusion equation.
We are now ready to state the long-time asymptotics. We obtain this result combining the hypocoercivity result for (1.2) provided by Theorem 2.2 and the exponential relaxation to equilibrium of the solution to the parabolic equation (1.3) on the torus due to Poincaré inequality.
@@
$\frac{1}{2}\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|\tilde{\rho}-1\|_{L_{x}^{2}}^{2}=-\int_{\mathbb{T}^{d}}\tilde{\rho}^{-\alpha}|\nabla_{x}\tilde{\rho}|^{2}\,\mathrm{d}x\leq-C\|\nabla_{x}\tilde{\rho}\|_{L_{x}^{2}}^{2}=-C\|\nabla_{x}(\tilde{\rho}-1)\|_{L_{x}^{2}}^{2}\leq\tilde{C}\|\tilde{\rho}-1\|^{2}_{L^{2}_{x}},$
@@ 5.2 Finite-time asymptotics
Let $\rho_{\mathrm{in}}\in C^{2}(\mathbb{T}^{d})$ and $h_{\mathrm{in}}\in L^{1}_{x,v}(\mathbb{T}^{d}\times\mathbb{R}^{d})$ both be valued in $[A^{-1},A]$ for $A>1$, and take $\varepsilon\in(0,\min\{\varepsilon_{0},1\})$ with $\varepsilon_{0}$ given by Theorem 2.2. Let $h_{\varepsilon}$ be the solution of (5.1) with initial data $h_{\mathrm{in}}$ and $\tilde{\rho}$ the solution of (1.3) with initial data $\rho_{\mathrm{in}}$. Then, there exist constants $C,C^{\prime}>0$ depending only on $\alpha,A,\|\rho_{\mathrm{in}}\|_{C^{2}}$ such that for all $t\geq 0$, the following estimate holds,
@@
Now we expect formally that as $\varepsilon\rightarrow 0$ we will have
@@
This means that (5.3) will hold in $H^{-1}_{t,x}([0,T]\times\mathbb{T}^{d})$ as it follows from the convergence of $j_{\varepsilon},E_{\varepsilon}$ to zero in $L^{2}([0,T]\times\mathbb{T}^{d})$. We aim to exploit this fact by rearranging the dissipation of the relative entropy. First, we complete the square with the two negative dissipation terms. To this end, we define
$Q_{\varepsilon}:=\frac{1}{\varepsilon}\rho_{\varepsilon}^{\frac{\alpha}{2}}j_{\varepsilon}+\tilde{\rho}^{-\frac{\alpha}{2}}\nabla_{x}\tilde{\rho},\quad\text{and}\quad R_{\varepsilon}:=\nabla_{x}\rho_{\varepsilon}+\varepsilon^{-1}\rho_{\varepsilon}^{\alpha}j_{\varepsilon}$
@@
$\int_{0}^{t}\int_{\mathbb{T}^{d}}e^{-Cs}R_{\varepsilon}\cdot\tilde{\rho}^{-\alpha}\nabla_{x}\tilde{\rho}\,\mathrm{d}s\,\mathrm{d}x$ $=\int_{0}^{t}\int_{\mathbb{T}^{d}}\Bigl{(}e^{-Cs}E_{\varepsilon}:\nabla_{x}\bigl{(}\tilde{\rho}^{-\alpha}\nabla_{x}\tilde{\rho}\bigr{)}+\varepsilon j_{\varepsilon}\cdot\partial_{s}\bigl{(}e^{-Cs}\tilde{\rho}^{-\alpha}\nabla_{x}\tilde{\rho}\bigr{)}\Bigr{)}\,\mathrm{d}s\,\mathrm{d}x$
@@
With this result, we are ready to show the global in time diffusive asymptotics.

### v3

@@ 4.1 $L^{2}$-hypocoercivity à la Dolbeault-Mouhot-Schmeiser
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\int_{\mathbb{T}^{d}}\int_{\mathbb{R}^{d}}\frac{(f_{\varepsilon}-\mathcal{M})^{2}}{\mathcal{M}}\,\mathrm{d}v\,\mathrm{d}x$ $=-\frac{1}{\varepsilon^{2}}\int_{\mathbb{T}^{d}}\rho_{\varepsilon}^{\alpha}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left(g_{\varepsilon}^{2}(u)\frac{\mathcal{M}(v)}{\mathcal{M}(u)}+g_{\varepsilon}^{2}(v)\frac{\mathcal{M}(u)}{\mathcal{M}(v)}-2g_{\varepsilon}(v)g_{\varepsilon}(u)\right)\,\mathrm{d}u\,\mathrm{d}v\,\mathrm{d}x$
$=-\frac{2}{\varepsilon^{2}}\int_{\mathbb{T}^{d}}\rho_{\varepsilon}^{\alpha}\int_{\mathbb{R}^{d}}\frac{g_{\varepsilon}^{2}}{\mathcal{M}}\,\mathrm{d}v\,\mathrm{d}x$
@@
Now using Lemma 4.1 and Young’s inequality we estimate the above quantity further as,
@@
$\qquad+\frac{1}{\varepsilon^{2}}\|\rho_{\varepsilon}^{\alpha}\|_{L^{\infty}}\left(\frac{\|\rho_{\varepsilon}^{\alpha}\|_{L^{\infty}}}{\varepsilon}\|j_{\varepsilon}\|_{L^{2}_{x}}^{2}+\frac{\varepsilon}{4\|\rho_{\varepsilon}^{\alpha}\|_{L^{\infty}}}\|\rho_{\varepsilon}-\bar{\rho_{\varepsilon}}\|_{L^{2}_{x}}^{2}\right)-\frac{C}{\varepsilon}\|\rho_{\varepsilon}-\bar{\rho_{\varepsilon}}\|_{L^{2}_{x}}^{2}$
$=-\frac{1}{\varepsilon}\bigl(C-\frac{1}{2}\bigl)\|\rho_{\varepsilon}-\bar{\rho_{\varepsilon}}\|_{L^{2}_{x}}^{2}+\frac{1}{\varepsilon}\|E_{\varepsilon}\|_{L^{2}_{x}}^{2}+\frac{1}{\varepsilon}\|j_{\varepsilon}\|_{L^{2}_{x}}^{2}+\frac{1}{\varepsilon^{3}}\|\rho_{\varepsilon}^{\alpha}\|_{L^{\infty}}^{2}\|j_{\varepsilon}\|_{L^{2}_{x}}^{2}$
$\leq-\frac{1}{4\varepsilon}\|\rho_{\varepsilon}-\bar{\rho_{\varepsilon}}\|_{L^{2}_{x}}^{2}+\frac{1}{\varepsilon^{3}}(2+\|\rho_{\varepsilon}^{\alpha}\|_{L^{\infty}}^{2})\|(I-\Pi)f_{\varepsilon}\|^{2}_{L^{2}_{x,v}(\mathcal{M}^{-1})},$ (4.5)
for $\varepsilon\leq 1$ where the last inequality follows from applying Lemma 2.5 with $\beta=\rho_{\varepsilon}$ and then integrating in space. To ease notation, we also bounded the constant appearing in Lemma 4.1 from below, $C=4\pi^{2}/(1+4\pi^{2})>3/4$.
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\left(\|f_{\varepsilon}-\mathcal{M}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}+\frac{\varepsilon}{A^{\alpha}(2+A^{2\alpha})}\int_{\mathbb{T}^{d}}j_{\varepsilon}\cdot(I-\Delta_{x})^{-1}\nabla_{x}\rho_{\varepsilon}\,\mathrm{d}x\right)\\ \leq-\frac{1}{\varepsilon^{2}}\frac{1}{A^{\alpha}}\|(I-\Pi)f_{\varepsilon}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}-\frac{1}{4A^{\alpha}(2+A^{2\alpha})}\|\rho_{\varepsilon}-\bar{\rho_{\varepsilon}}\|_{L^{2}_{x}}^{2}.$ (4.6)
@@
$\frac{\varepsilon}{A^{\alpha}(2+A^{2\alpha})}\int_{\mathbb{T}^{d}}j_{\varepsilon}\cdot(I-\Delta_{x})^{-1}\nabla_{x}\rho_{\varepsilon}\,\mathrm{d}x$ $\leq\frac{\varepsilon}{A^{\alpha}(2+A^{2\alpha})}\|j_{\varepsilon}\|_{L^{2}_{x}}\|(I-\Delta_{x})^{-1}\nabla_{x}\rho_{\varepsilon}\|_{L^{2}_{x}}$
$\leq\frac{\varepsilon}{A^{\alpha}(2+A^{2\alpha})}\|f_{\varepsilon}-\mathcal{M}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}\leq\frac{1}{2}\|f_{\varepsilon}-\mathcal{M}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})},$
with the second equality following from Lemma 2.5 with $\beta=1$. Consequently for $\varepsilon$ sufficiently small, we obtain
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\Big($ $\|f_{\varepsilon}-\mathcal{M}\|^{2}_{L^{2}(\mathcal{M}^{-1})}+\frac{\varepsilon}{A^{\alpha}(2+A^{2\alpha})}\int_{\mathbb{T}^{d}}j_{\varepsilon}\cdot(I-\Delta_{x})^{-1}\nabla_{x}\rho_{\varepsilon}\,\mathrm{d}x\Big)$
$\leq-\frac{1}{4A^{\alpha}(2+A^{2\alpha})}\bigl(\|(I-\Pi)f_{\varepsilon}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}+\|\rho_{\varepsilon}-\bar{\rho_{\varepsilon}}\|_{L^{2}_{x}}^{2}\bigr)$
$=-\frac{1}{4A^{\alpha}(2+A^{2\alpha})}\|f_{\varepsilon}-\mathcal{M}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}$
$\leq-\frac{1}{8A^{\alpha}(2+A^{2\alpha})}\left(\|f_{\varepsilon}-\mathcal{M}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}+\frac{\varepsilon}{A^{\alpha}(2+A^{2\alpha})}\int_{\mathbb{T}^{d}}j_{\varepsilon}\cdot(I-\Delta_{x})^{-1}\nabla_{x}\rho_{\varepsilon}\,\mathrm{d}x\right).$
@@
$\|f_{\varepsilon}(t)-\mathcal{M}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}+\frac{\varepsilon}{A^{\alpha}(2+A^{2\alpha})}\int_{\mathbb{T}^{d}}j_{\varepsilon}(t)\cdot(I-\Delta_{x})^{-1}\nabla_{x}\rho_{\varepsilon}(t)\,\mathrm{d}x\\ \leq e^{-\gamma t}\left(\|f_{\varepsilon}(0)-\mathcal{M}\|^{2}_{L_{x,v}^{2}(\mathcal{M}^{-1})}+\frac{\varepsilon}{A^{\alpha}(2+A^{2\alpha})}\int_{\mathbb{T}^{d}}j_{\varepsilon}(0)\cdot(I-\Delta_{x})^{-1}\nabla_{x}\rho_{\varepsilon}(0)\,\mathrm{d}x\right),$
@@ 5.1 Long-time asymptotics
Notice that Equation (1.3) is uniformly parabolic away from zero and $\infty$, so the classical theory applies for initial data uniformly bounded from above and below. We then have well-posedness and a strong maximum principle and Lemma 5.1 follows. We refer the reader to [38] for a review of the theory of the nonlinear diffusion equation.
We are now ready to state the long-time asymptotics. We obtain this result combining the hypocoercivity result for (1.2) provided by Theorem 2.2 and the exponential relaxation to equilibrium of the solution to the parabolic equation (1.3) on the torus due to the Poincaré inequality.
@@
$\frac{1}{2}\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|\tilde{\rho}-1\|_{L_{x}^{2}}^{2}=\int_{\mathbb{T}^{d}}(\tilde{\rho}-1)\partial_{t}\tilde{\rho}\,\mathrm{d}x=\int_{\mathbb{T}^{d}}\tilde{\rho}\partial_{t}\tilde{\rho}\,\mathrm{d}x=\int_{\mathbb{T}^{d}}\tilde{\rho}\nabla_{x}\cdot\bigl(\tilde{\rho}^{-\alpha}\nabla_{x}\tilde{\rho}\bigr)\,\mathrm{d}x.$
where the second inequality follows from the conservation of mass. Integrating by parts, we then find
$\frac{1}{2}\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|\tilde{\rho}-1\|_{L_{x}^{2}}^{2}=-\int_{\mathbb{T}^{d}}\tilde{\rho}^{-\alpha}|\nabla_{x}\tilde{\rho}|^{2}\,\mathrm{d}x\leq-C\|\nabla_{x}\tilde{\rho}\|_{L_{x}^{2}}^{2}=-C\|\nabla_{x}(\tilde{\rho}-1)\|_{L_{x}^{2}}^{2}\leq-\tilde{C}\|\tilde{\rho}-1\|^{2}_{L^{2}_{x}},$
@@ 5.2 Finite-time asymptotics
Let $\rho_{\mathrm{in}}\in C^{2}(\mathbb{T}^{d})$ and $h_{\mathrm{in}}\in L^{1}_{x,v}(\mathbb{T}^{d}\times\mathbb{R}^{d})$ both be valued in $A^{-1},A]$ for $A>1$, and take $\varepsilon\in(0,\min\{\varepsilon_{0},1\})$ with $\varepsilon_{0}$ given by Theorem 2.2. Let $h_{\varepsilon}$ be the solution of (5.1) with initial data $h_{\mathrm{in}}$ and $\tilde{\rho}$ the solution of (1.3) with initial data $\rho_{\mathrm{in}}$. Then, there exist constants $C,C^{\prime}>0$ depending only on $\alpha,A,\|\rho_{\mathrm{in}}\|_{C^{2}}$ such that for all $t\geq 0$, the following estimate holds,
@@
Now, by Remark (2.8), we expect formally that as $\varepsilon\rightarrow 0$ we will have
@@
This means that (5.3) does in fact hold in $H^{-1}_{t,x}([0,T]\times\mathbb{T}^{d})$ as it follows from the convergence of $j_{\varepsilon},E_{\varepsilon}$ to zero in $L^{2}([0,T]\times\mathbb{T}^{d})$. We give a detailed proof of this later on. We aim to exploit this fact by rearranging the dissipation of the relative entropy. First, we complete the square with the two negative dissipation terms. To this end, we define
$Q_{\varepsilon}:=\frac{1}{\varepsilon}\rho_{\varepsilon}^{\frac{\alpha}{2}}j_{\varepsilon}+\tilde{\rho}^{-\frac{\alpha}{2}}\nabla_{x}\tilde{\rho},\quad\text{and}\quad R_{\varepsilon}:=\nabla_{x}\rho_{\varepsilon}+\frac{1}{\varepsilon}\rho_{\varepsilon}^{\alpha}j_{\varepsilon}$
@@
$\int_{0}^{t}\int_{\mathbb{T}^{d}}e^{-Cs}R_{\varepsilon}\cdot\tilde{\rho}^{-\alpha}\nabla_{x}\tilde{\rho}\,\mathrm{d}x\,\mathrm{d}s$ $=\int_{0}^{t}\int_{\mathbb{T}^{d}}\Bigl(e^{-Cs}E_{\varepsilon}:\nabla_{x}\bigl(\tilde{\rho}^{-\alpha}\nabla_{x}\tilde{\rho}\bigr)+\varepsilon j_{\varepsilon}\cdot\partial_{s}\bigl(e^{-Cs}\tilde{\rho}^{-\alpha}\nabla_{x}\tilde{\rho}\bigr)\Bigr)\,\mathrm{d}x\,\mathrm{d}s$
@@
With this result, we are ready to show the global in time diffusive asymptotics. Recall that in the following we take $0<\varepsilon,\varepsilon^{\prime}<\frac{1}{2}$, where $\varepsilon$ denotes our scaling parameter and $\varepsilon^{\prime}$ is a bound on the distance between the initial data for the kinetic and parabolic equations, i.e. $\|f_{\varepsilon,\mathrm{in}}-\rho_{\mathrm{in}}\mathcal{M}\|_{L_{x,v}^{2}(\mathcal{M}^{-1})}\leq\varepsilon^{\prime}$.